In a triangle ABC, angle B's measurement is 3 times angle A's measurement, and angle C's measurement is 4 degrees less than 6 times angle A's measurement. If a triangle measures 180 degrees all together, what is the measurement of angle A?
Answer by mananth(16949)

Let angle A be x
Angle B will be 3x
Angle C = 6x-4
x+3x+6x-4=180
10x=180-4
10x=176
x= 176/10
x=17.6 degrees